Solve
log2(x-9)=4

Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(32388)
Answer by ewatrrr(24785)


Hi,
Solve
log2(x-9)=4

2^4 = (x-9)
16 = x - 9
25 = x